OpenStudy (anonymous):

Find the product. 2 5/6 * 6 4/7= Enter answer as mixed number in simplest terms.

OpenStudy (mertsj):

Can you change those mixed numbers to improper fractions?

OpenStudy (anonymous):

i don't know.

OpenStudy (mertsj):

Well why don't you try and tell me what you get.

OpenStudy (anonymous):

I dont even know where to begin.

OpenStudy (anonymous):

idk how to do that junk.

OpenStudy (mertsj):

I'll do the first one and then you see if you can get the second one.

OpenStudy (mertsj):

\[2 \frac{5}{6}=\frac{6\times 2 + 5}{6}=\frac{17}{6}\]

OpenStudy (mertsj):

Now you try 6 and 4/7

OpenStudy (mertsj):

\[6 \frac{4}{7}=\]

OpenStudy (anonymous):

7x4+6/7= ummm idk.......

OpenStudy (anonymous):

or 7x6+4/7

OpenStudy (mertsj):

Multiply the 6 and 7 What is 6 times 7?

OpenStudy (anonymous):

42+4=46

OpenStudy (anonymous):

46/7

OpenStudy (mertsj):

Good. Put that over 7

OpenStudy (anonymous):

i did

OpenStudy (mertsj):

\[\frac{17}{6}\times \frac{46}{7}\]

OpenStudy (anonymous):

ummm

OpenStudy (mertsj):

Notice that 6 and 46 have the common factor of 2. Cancel that out.

OpenStudy (mertsj):